# Jarne Verhaeghe IDLab

### v1

@@ 4.1.3 Conformal Predictive Systems
In some cases, providing a prediction interval often does not suffice and a complete predictive distribution is required. The extension proposed by Vovk et al. (2019) produces a predictive distribution by arranging p-values, created using specific conformity measures, into a probability distribution function. A requirement to create a Conformal Predictive System (CPS) is to use a specific type of conformity measures  [^1] For the specific definition see  Vovk et al. (2020) which include monotonic measures. Then, given the training data $Z_{1:n}$ and observed test sample $X_{n+1}$, we define an example of this specific conformity measure $S$ and conformity scores $R_{i}^{y}$ similar as in equations 3 and 4:
@@
With $\hat{\mu}$ an estimator fitted on the training set $Z_{1:n}$. $R_{i}^{y}$ and $R_{n+1}^{y}$ are then similarly defined as in equation 3 for a CPS. Then, as defined in Vovk et al. (2022) we can define a predictive distribution Q for value $y$, using a distribution of nonconformity scores $F_{R}(y)$ of $y$ to calculate $\mathbb{P}$, similarly to the quantile function in equation 6 as follows:
@@
Where $\phi$ is a random number sampled from a uniform distribution between 0 and 1 to ensure a smooth predictive distribution. Using the same approach as section 4.1.2, these conformal predictive systems can be expanded to weighted conformal predictive systems by adjusting $F_{R}(y)$ to account for the covariate shift (Jonkers et al., 2024a).
Additionally, conformal predictive systems also suffer from computational issues, therefore Vovk et al. (2020) introduced split conformal predictive systems to tackle the same issues in a way analogous to section 4.1.1.
@@ 4.2 Proposed methodology: Propensity Weighted Conformal Prediction
Taking into account the background knowledge of conformal prediction, we first need to formally define the target distribution to tackle our problem definition. A CADRF model $\hat{\nu}(X,T)$ is trained on triples $(X,T,Y)$ with X $d$-dimensional observed covariates $X\in\mathbb{R}^{d}\sim P_{X}$ and continuous treatment variables $T\in[t_{L},t_{U}]\sim P_{T|X}$ to predict responses $Y\in\mathbb{R}\sim P_{Y|T,X}$. $P_{X}$ represents the covariate distribution, $P_{T|X}$ represents the observational conditional treatment distribution given confounders $X$, and $P_{Y|T,X}$ represents the outcome distribution. $P_{T|X}=P_{T}$ if there are no confounders for T. A CADRF model will be used to query the dose-response for all $T\in[t_{L},t_{U}]$, creating an interventional distribution $\tilde{P}_{T}$. As every treatment value $t$ is equally likely we can define $\tilde{P}_{T}=\tilde{P}_{T|X}=Uniform(t_{L},t_{U})$.
To attain marginal coverage across the interventional test set for a CADRF we can use weighted conformal prediction (Tibshirani et al., 2019). This requires defining the weights $w$ for $X_{i}$ and treatment value $t$ using equation 11, which we will call the global ($g$) propensity ($p$) weights $w_{g,p}$:
@@
$=\frac{d\tilde{P}_{T|X}(X_{i},T_{i})}{dP_{T|X}(X_{i},T_{i})}=\frac{f_{U(t_{L},t_{U})}(T_{i})}{\pi(T_{i}|X_{i})}=\frac{\frac{\mathbbm{1}_{[t_{L},t_{U}]}(T_{i})}{t_{U}-t_{L}}}{\pi(T_{i}|X_{i})}\propto\frac{\mathbbm{1}_{[t_{L},t_{U}]}(T_{i})}{\pi(T_{i}|X_{i})}$
with $\mathbbm{1}_{[t_{L},t_{U}]}(T_{i})$ the indicator function for $T_{i}\in[t_{L},t_{U}]$.
We assume that there is no distribution shift for $X$ and thus $\tilde{P}_{X}(X_{i})=P_{X}(X_{i})$. Additionally, $f_{U(t_{L},t_{U})}$ is the probability density function for the uniform distribution. We also define the propensity function $\pi(T_{i}|X_{i})$ as the probability density function for $P_{T|X}(T_{i})$ as specified in Section 2. To generate the prediction intervals at treatment value $t$ for a new sample $X_{n+1}$ the weights change to $w_{g,p}(X_{n+1},t)=\frac{1}{\pi(t|X_{n+1})}$. According to the weighted exchangeability defined in (Tibshirani et al., 2019), this guarantees marginal coverage over the interventional distribution, for all $T\in[t_{L},t_{U}]$, and $X\sim P_{X}$.
Tibshirani et al. (2019) also suggested a method to attain local coverage around a predetermined target point $x_{0}$ using weighted conformal prediction. Consequently, this can provide varying prediction intervals for different values of $x_{0}$ providing another heteroscedastic approach. The proposed weights, which we call the local ($l$) weights $w_{l}$, utilize kernel functions with bandwidth parameter $h$:
@@
These weights then guarantee
@@
This assures coverage around $x_{0}$, but $x_{0}$ must be determined beforehand. Additionally, if a new $x_{0}$ must be evaluated, a new calibration procedure must be performed which should be considered when applying it to general regression use cases. However, for this work, the target interventional treatment distribution is known in advance and can all be computed before deployment. Consequently, for a target treatment value $t$ we can define $w_{l}^{t}(T_{i})\propto K(\frac{T_{i}-t}{h})$ instead.
The local weights guarantee coverage where $d\tilde{P}_{T}(T_{i})/dP_{T}(T_{i})\propto K(\frac{T_{i}-t}{h})$. To adjust the local weights for a CADRF model we need to be aware of the covariate shift introduced by evaluating the interventional distribution and thus must combine $w_{g,p}$ with $w_{local}$ to achieve weighted exchangeability. These new weights are defined as $w_{l,p}$ for target treatment $t$:
$w_{l,p}^{t}(X_{i},T_{i})\propto\frac{\mathbbm{1}_{[t_{L},t_{U}]}(T_{i})K\left(\frac{T_{i}-t}{h}\right)}{\pi(T_{i}|X_{i})}$ (18)
To generate the prediction intervals for target treatment $t$ for a new sample $X_{n+1}$ the weights are then $w_{l,p}^{t}(X_{n+1},t)=\frac{\mathbbm{1}_{[t_{L},t_{U}]}(T_{i})K((t-t/h))}{\pi(t|X_{i})}=\frac{\mathbbm{1}_{[t_{L},t_{U}]}(T_{i})}{\pi(t|X_{i})}$, which is equal to $w_{g,p}^{t}(X_{n+1},t)$. By using these weights in a weighted conformal prediction framework, we provide a solution to the problem definition in Section 2.
@@ 5.1 Synthetic Data
We evaluate the proposed approach on synthetic data as evaluating the true individual dose-response curve requires knowing the counterfactuals which is not feasible in real-world data.
We used three experimental setups using synthetic data, each having different scenarios that change specific parameters. Setup 1 is inspired by Wu et al. (2024) and Setup 2 follows the experimental setup of Schröder et al. (2024). Both Setup 1 and 2 are clarified in Appendix A. Setup 3 is novel, proposed by us, which mimics a situation where, for every scenario, two different possible dose-response functions are possible that each depends on the covariates, resulting in heavy confounding and thus limited overlap.
For each scenario (over the different setups), 5000 samples were generated using 50 different random seeds resulting in 50 datasets for each scenario. These datasets were split into 25% test (1250), 25% calibration (1250), and 50% training (2500) samples. For each scenario, two different $\alpha$ (significance values) were evaluated (i.e., $0.1$ and $0.05$ for a confidence of $90\%$ and $95\%$ resp.). Each sample in the test set is evaluated using 40 treatment values $t_{0}$ at equal intervals between the $2\%$ and $98\%$ training treatment value quantile to include varying treatment overlap regions and to mimic the uniform treatment sampling. In the results, the coverage of all treatment values and all samples in the test set are aggregated to a single mean coverage for each experiment, resulting in 50 mean coverage results for every method and scenario.
@@ 5.1.1 Setup 3
The two scenarios have slightly different outcome distributions, as shown in Table 1. The idea is the same for both scenarios; The individual dose-response function is truly conditional and thus equal treatment values between different individuals or samples do not necessarily translate to each other. In total, there are four different possible dose-response functions depending on the covariates. Furthermore, there is heavy confounding resulting in limited samples where $T-X_{2}$ yields high values that in turn create large outcome values. This creates an opportunity for high epistemic uncertainty and limited overlap. For scenario two, the aleatoric uncertainty is also heteroscedastic based on $X_{3}$ forcing solutions to look beyond the treatment value to quantify uncertainty.
@@
$1$ $Y\sim sign(X_{3})\cdot(2(T-X_{2}))^{2}+33T\cdot sign(X_{1})+\text{Normal}(0,2)$
$2$ $Y\sim sign(X_{3})\cdot(2(T-X_{2}))^{2}+33T\cdot sign(X_{1})$
$+\frac{(sign(X_{3})+1)}{2}\cdot\text{Normal}(0,30)+\text{Normal}(0,2)$

### v2

@@ 4.1.3 Conformal Predictive Systems
In some cases, providing a prediction interval often does not suffice and a complete predictive distribution is required. The extension proposed by vovk_nonparametric_2019 produces a predictive distribution by arranging p-values, created using specific conformity measures, into a probability distribution function. A requirement to create a Conformal Predictive System (CPS) is to use a specific type of conformity measures  [^1] For the specific definition see  vovk_computationally_2020 which include monotonic measures. Then, given the training data $Z_{1:n}$ and observed test sample $X_{n+1}$, we define an example of this specific conformity measure $S$ and conformity scores $R_{i}^{y}$ similar as in equations 3 and 4:
@@
With $\hat{\mu}$ an estimator fitted on the training set $Z_{1:n}$. $R_{i}^{y}$ and $R_{n+1}^{y}$ are then similarly defined as in equation 3 for a CPS. Then, as defined in vovk_algorithmic_2022 we can define a predictive distribution $Q$ for value $y$, using a distribution of nonconformity scores $F_{R}(y)$ of $y$ to calculate $\mathbb{P}$, similarly to how the quantile function in equation 6 for standard conformal prediction was defined:
@@
Where $\phi$ is a random number sampled from a uniform distribution between 0 and 1 to ensure a smooth predictive distribution. Using the same approach as section 4.1.2, these conformal predictive systems can be expanded to weighted conformal predictive systems by adjusting $F_{R}(y)$ to account for the covariate shift (jonkers_conformal_pred_shift_2024).
Additionally, conformal predictive systems also suffer from computational issues, therefore vovk_computationally_2020 introduced split conformal predictive systems to tackle the same issues in a way analogous to section 4.1.1.
@@ 4.2 Proposed methodology: Propensity Weighted Conformal Prediction
Taking into account the background knowledge of conformal prediction, we first need to formally define the target distribution to tackle our problem definition. A CADRF model $\hat{\nu}(X,T)$ is trained on triples $(X,T,Y)$ with X $d$-dimensional observed covariates $X\in\mathbb{R}^{d}\sim P_{X}$ and continuous treatment variables $T\in[t_{L},t_{U}]\sim P_{T|X}$ to predict responses $Y\in\mathbb{R}\sim P_{Y|T,X}$. $P_{X}$ represents the covariate distribution, $P_{T|X}$ represents the observational conditional treatment distribution given confounders $X$, and $P_{Y|T,X}$ represents the outcome distribution. $P_{T|X}=P_{T}$ if there are no confounders for T. The CADRF model will be used to query the dose-response for all $T\in[t_{L},t_{U}]$. This query simulates an intervention creating an interventional distribution $\tilde{P}_{T}$ where $\tilde{P}_{T|X}=\tilde{P}_{T}$, because we query $T$, irregardless of X. In the case that every treatment value $t$ is equally likely to be queried, we can define this interventional distribution as: $\tilde{P}_{T}=Uniform(t_{L},t_{U})$.
We can utilize this uniform interventional distribution to ensure that the model is calibrated for every possible dose equally across the domain. This guarantees that the model can be queried to provide a result for any specific dose $t$ within the defined range. It is crucial to distinguish this methodological choice (uniform intervention) from the observational treatment distribution ($P_{T|X}$), which reflects how doses actually occur in a real-world setting in observational data. The core idea here is that we are modeling an interventional distribution ($\tilde{P}_{T}$) where every dose $t$ is equally likely to be queried, and not the underlying observational distribution ($P_{T|X}$) from which the data was collected. However, the results of this paper are not limited to this uniform interventional distribution. If a use-case requires a known non-uniform interventional distribution, substituting the $\text{Uniform}(t_{L},t_{U})$ distribution with the appropriate $\tilde{P}_{T}$ will still yield valid results. Following, all derivations will be performed using a general interventional distribution $\tilde{P}_{T}$ with probability density function $f_{\tilde{P}_{T}}(T)$ with the uniform interventional distribution as an example.
To attain marginal coverage across the interventional test set for a CADRF we can use weighted conformal prediction (tibshirani_conformal_2019). This requires defining the weights $w$ for $X_{i}$ and treatment value $t$ using equation 11, which we will call the global ($g$) propensity ($p$) weights $w_{g,p}$:
@@
$=\frac{d\tilde{P}_{T|X}(X_{i},T_{i})}{dP_{T|X}(X_{i},T_{i})}=\frac{d\tilde{P}_{T}(X_{i},T_{i})}{dP_{T|X}(X_{i},T_{i})}=\frac{f_{\tilde{P}_{T}}(T_{i})}{\pi(T_{i}|X_{i})}$
Now, in the case of a uniform interventional distribution $\text{Uniform}(t_{L},t_{U})$ we get the following global weights:
$w_{g,p,uniform}(X_{i},T_{i})$ $=\frac{f_{\tilde{P}_{T}}(T_{i})}{\pi(T_{i}|X_{i})}=\frac{\frac{\mathbbm{1}_{[t_{L},t_{U}]}(T_{i})}{t_{U}-t_{L}}}{\pi(T_{i}|X_{i})}\propto\frac{\mathbbm{1}_{[t_{L},t_{U}]}(T_{i})}{\pi(T_{i}|X_{i})}$ (16)
with $\mathbbm{1}_{[t_{L},t_{U}]}(T_{i})$ the indicator function for $T_{i}\in[t_{L},t_{U}]$ and $f_{U(t_{L},t_{U})}$ the probability density function for the uniform distribution.
For simplicity, we assume that there is no distribution shift for $X$ and thus $\tilde{P}_{X}(X_{i})=P_{X}(X_{i})$. We also define the propensity function $\pi(T_{i}|X_{i})$ as the probability density function for $P_{T|X}(T_{i})$ as specified in Section 2. To generate the prediction intervals at treatment value $t$ for a new sample $X_{n+1}$ the weights change to $w_{g,p}(X_{n+1},t)=\frac{1}{\pi(t|X_{n+1})}$. According to the weighted exchangeability defined in (tibshirani_conformal_2019), this guarantees marginal coverage over the interventional distribution, for all $T\in[t_{L},t_{U}]$, and $X\sim P_{X}$. tibshirani_conformal_2019 also suggested a method to attain local coverage around a predetermined target point $x_{0}$ using weighted conformal prediction. Consequently, this can provide varying prediction intervals for different values of $x_{0}$, providing another heteroscedastic approach. The proposed weights, which we call the local ($l$) weights $w_{l}$, utilize kernel functions with bandwidth parameter $h$:
@@
These weights then guarantee tibshirani_conformal_2019
@@
The bandwidth $h$ is a hyperparameter whose choice directly impacts the efficiency of the prediction interval. To select $h$, we advise evaluating the calibration process to set $h$ small enough while having enough samples in the neighbourhood of $x_{0}$. Additionally, $x_{0}$ must be determined beforehand. If a new $x_{0}$ must be evaluated, a new calibration procedure must be performed which should be considered when applying it to general regression use cases. However, for this work, the target interventional treatment distribution is known in advance and can all be computed before deployment. For the target interventional treatment distribution, we can translate $x_{0}$ to a target treatment value $t$ and define $w_{l}^{t}(T_{i})\propto K(\frac{T_{i}-t}{h})$ instead. The local weights guarantee coverage where $d\tilde{P}_{T}(T_{i})/dP_{T}(T_{i})\propto K(\frac{T_{i}-t}{h})$. The coverage guarantee for $t$, similar to 18, then directly follows from Proposition 1. The theoretical proofs for this proposition are in Appendix A.
Proposition 1 (following tibshirani_conformal_2019; lei_conformal_2021).
Assume $(X_{i},T_{i},Y_{i})\overset{i.i.d.}{\sim}P_{X}\times P_{T|X}\times P_{Y|T,X}$, i=1,,ni=1,...,n; the likelihood ratio $w(X,T)\propto\frac{d\tilde{P}_{T|X}}{dP_{T|X}}$; and the estimated likelihood ratio $\hat{w}(X,T)$. Using WCP to construct $\hat{C}(X,T)$, the following finite-sample bounds apply:
  • S1.
    (Oracle Likelihood Ratio) If $\hat{w}(\cdot,\cdot)=w(\cdot,\cdot)$, i.e. oracle likelihood ratio function; then,
    $1-\alpha\leq\mathbb{P}_{(X,T,Y)\sim P_{X}\times\tilde{P}_{T|X}\times P_{Y|T,X}}\{Y\in\hat{C}(X,T)\}$ (19)
  • S2.
    (Finite Sample with Regularity Conditions) If $\hat{w}(\cdot,\cdot)=w(\cdot,\cdot)$; the non-conformity scores $S_{i}$ have no ties almost surely; $\tilde{P}_{T|X}\times P_{X}$ is absolutely continuous with respect to $P_{T|X}\times P_{X}$; and $(\mathbb{E}_{(X,T)\sim P_{X}\times P_{T|X}}[w(X,T)^{r}])^{\frac{1}{r}}\leq M_{r}<\infty$ where $r>0$ and $M_{r}$ denotes the upper bound of the $r$-th moment of the likelihood ratio; then,
    $1-\alpha\leq\mathbb{P}_{(X,T,Y)\sim P_{X}\times\tilde{P}_{T|X}\times P_{Y|T,X}}\{Y\in\hat{C}(X,T)\}\leq 1-\alpha+cn^{\frac{1}{r-1}}$ (20)
    where $c$ is an arbitrary positive constant depending on $M_{r}$ and $r$.
  • S3.
    (Estimated Likelihood Ratio) If $\hat{w}(\cdot,\cdot)\neq w(\cdot,\cdot)$; $\Delta_{w}=\frac{1}{2}\mathbb{E}_{(X,T)\sim P_{X}\times P_{T|X}}[|\hat{w}(X,T)-w(X,T)|]$; $(\mathbb{E}_{(X,T)\sim P_{X}\times P_{T|X}}[\hat{w}(X,T)^{r}])^{\frac{1}{r}}\leq M_{r}<\infty$; and further assuming the same assumptions as in S2.; then,
    $1-\alpha-\Delta_{w}\leq\mathbb{P}_{(X,T,Y)\sim P_{X}\times\tilde{P}_{T|X}\times P_{Y|T,X}}\{Y\in\hat{C}(X,T)\}\leq 1-\alpha+\Delta_{w}+cn^{\frac{1}{r-1}}$ (21)
To adjust the local weights for a CADRF model we need to be aware of the covariate shift introduced by evaluating the interventional distribution and thus must combine $w_{g,p}$ with $w_{local}$ to achieve weighted exchangeability. These new weights are defined as $w_{l,p}$ for target treatment $t$:
$w_{l,p}^{t}(X_{i},T_{i})\propto\frac{f_{\tilde{P}_{T}}(T_{i})K\left(\frac{T_{i}-t}{h}\right)}{\pi(T_{i}|X_{i})}$ (22)
Similarly to the global weights, in the case of a uniform interventional distribution $\text{Uniform}(t_{L},t_{U})$ we get the following local weights:
$w_{l,p,uniform}^{t}(X_{i},T_{i})\propto\frac{\mathbbm{1}_{[t_{L},t_{U}]}(T_{i})K\left(\frac{T_{i}-t}{h}\right)}{\pi(T_{i}|X_{i})}$ (23)
Once the calibration weights have been calculated and thus the calibration has been performed, we can infer the prediction intervals for target treatment $t$ and sample $X_{n+1}$ from these weights. To generate the prediction intervals for target treatment $t$ for a new sample $X_{n+1}$, we need the weight of the new sample $w_{l,p}^{t}(X_{n+1},t)$ for equation 11. Then, using equation 23 and $T_{n+1}=t$, this weight becomes $w_{l,p}^{t}(X_{n+1},t)=\frac{f_{\tilde{P}_{T}}(T_{i})K((t-t/h))}{\pi(t|X_{i})}=\frac{f_{\tilde{P}_{T}}(T_{i})}{\pi(t|X_{i})}$, which is equal to $w_{g,p}^{t}(X_{n+1},t)$. By using these weights in a weighted conformal prediction framework, we provide a solution to the problem definition in Section 2.
@@ 5.1 Synthetic Data
We evaluate the proposed approach on synthetic and semi-synthetic data as evaluating the true individual dose-response curve requires knowing the counterfactuals which is simply not possible in real-world data. Therefore, to evaluate the method we are forced to use (semi-)synthetic data. For the synthetic benchmarking, we used three experimental setups using synthetic data, each having different scenarios that change specific parameters. Setup 1 is inspired by wu_matching_2024 and Setup 2 follows the experimental setup of schroder_conformal_2024. Both Setup 1 and 2 are below. Setup 3 is novel, proposed by us, which mimics a situation where, for every scenario, two different possible dose-response functions are possible that each depends on the covariates, resulting in heavy confounding and thus limited overlap. For each scenario (over the different setups), 5000 samples were generated using 50 different random seeds resulting in 50 datasets for each scenario. These datasets were split into 25% test (1250), 25% calibration (1250), and 50% training (2500) samples. For each scenario, two different $\alpha$ (significance values) were evaluated (i.e., $0.1$ and $0.05$ for a confidence of $90\%$ and $95\%$ resp.). Each sample in the test set is evaluated using 40 treatment values $t_{0}$ at equal intervals between the $2\%$ and $98\%$ training treatment value quantile to include varying treatment overlap regions and to mimic the uniform treatment sampling. In the results, the coverage of all treatment values and all samples in the test set are aggregated to a single mean coverage for each experiment, resulting in 50 mean coverage results for every method and scenario.
5.2 Setup 1
For setup 1, inspired by wu_matching_2024, six independent covariates are sampled from various distributions representing both continuous and discrete values:
$X_{1},X_{2},X_{3},X_{4}$ $\sim\text{Normal}(0,1)$
$X_{5}$ $\sim\text{Uniform}[-2,2]\text{ (Integer)}$
$X_{6}$ $\sim\text{Uniform}(-3,3)$
Scenario Treatment function
33 $T_{\mu}=-0.8+X_{1}+0.1X_{2}-0.1X_{3}+0.2X_{4}+0.1X_{5}+0.1X_{6}+\frac{3}{2}X_{3}^{2}$
Table 1: The treatment functions for all scenarios in setup 1.
Scenario Treatment $T$ Treatment Assignment Distribution
1 $9T_{\mu}+17$ $T+\text{Normal}(0,5)$
2 $15T_{\mu}+22$ $T+\text{StudentT}(df=2)$
3 $9T_{\mu}+15$ $T+\text{Normal}(0,5)$
4 $49\frac{e^{T_{\mu}}}{1+e^{T_{\mu}}}-6$ $T+\text{Normal}(0,5)$
5 $42\frac{1}{1+e^{T_{\mu}}}+18$ $T+\text{Normal}(0,5)$
6 $7log(|T_{\mu}|+0.001)+13$ $T+\text{Normal}(0,4)$
7 $7T_{\mu}+16$ $T+\text{Normal}(0,1)$
8 $7T_{\mu}+16$ $T+20\cdot\text{Beta}(\alpha=2,\beta=8)$
Table 2: The propensity functions per scenario for Setup 1
$Y$ $\sim-1-(2X_{1}+2X_{2}+3X_{3}^{3}-20X_{4}-2X_{5}+20X_{6})$
5.3 Setup 2
Setup 2 tests the different treatment assignment distributions in the two different scenarios, which is the same experimental setup as proposed by schroder_conformal_2024. The covariates are sampled from a discrete uniform distribution. The treatment is sampled from the treatment assignment distributions shown in Table 3. The outcome function is sampled from a normal distribution with a mean determined by a sinus function based on both $X$ and $T$:
$X$ $\sim\text{Uniform}[1,4]\text{ (Integer)}$
$Y$ $\sim sin\left((0.05\pi)(T-X)\right)+\text{Normal}(0,0.1)$
Scenario Treatment Assignment Distribution
2 $T\sim\text{Normal}(5X,10)$
Table 3: The propensity functions per scenario for Setup 2
@@ 5.3.1 Setup 3
The two scenarios have slightly different outcome distributions, as shown in Table 4. The idea is the same for both scenarios; The individual dose-response function is conditional, and thus equal treatment values between different individuals or samples do not necessarily translate to each other. In total, there are four different possible dose-response functions depending on the covariates. Furthermore, there is heavy confounding resulting in limited samples where $T-X_{2}$ yields high values that in turn create large outcome values. This creates an opportunity for high epistemic uncertainty and limited overlap. For scenario two, the aleatoric uncertainty is also heteroscedastic based on $X_{3}$ forcing solutions to look beyond the treatment value to quantify uncertainty.
@@
11 $Y\sim sign(X_{3})\cdot(2(T-X_{2}))^{2}+33T\cdot sign(X_{1})+\text{Normal}(0,2)$
22 $Y\sim sign(X_{3})\cdot(2(T-X_{2}))^{2}+33T\cdot sign(X_{1})$
$\frac{(sign(X_{3})+1)}{2}\cdot\text{Normal}(0,30)+\text{Normal}(0,2)$